1. find the range of f(x)-2x^2+4 for the domain (-3,0,2)
2.given the arithmetic sequence A(n)=5+(n-1)d. Find the first and third terms if the common difference is 2.
3.given a Geometric sequence, A(n)=A=(1)r^n-1. Find the following terms if the first term is 3 and the common ratio is -2. Find the first, second and third term.
4. What is the slope of the line through the points (-1,2) and (2,-3).
5. A=1/2h(x+); solve for x.
6. Which of these equations is a direct variation? What is the constant of variation for those that are a direst variation? (a) y=x/2,(b)3x-5y=0 (c)y=4-x
7. Tell whether the lines for the pair of equations are parallel, perpendicular or neither: y=(-1/3)x-5, -18x+6y=21
8. Write the equation of a line in slope-intercept form that is perpendicular to the given line and that passes through the given point: y=(2/5)x+6;(-4,8)
9. find the x and y intercepts of the line 7x+6y=14

Answer by edjones(8007)

9)
7x+6y=14
6y=14-7x
y = 7/3 - 7x/6
.
7/3 - 7x/6 = 0
14-7x=0 multiply each side by 6.
-7x=-14
x=2 the x intercept.
.
y=7/3 the y intecept when x=0.